Given a rectangular prism with dimensions w = 3, l = 4, and h = 6. If you created a second rectangular prism with the length doubled but the height halved (and the width stays the same), which would be the relation of the second volume to the first volume?
Answer by drk(1908)

The volume of our first prism is
V = LWH
or
V = 3*4*6 = 72
--
The volume of the second prism is
V = LWH
however, we change the information and from above, we get
V = (2L)(.5H)W
which is
V = (2*3)(.5*4)(6) = 6*2*6 = 72
--
The volumes are still the same.